One number is 5 times as large as another. The sum of their reciprocals is 3/5. What are the numbers?
Answer by htmentor(1343)

Let x = the smaller number
Then the larger number = 5x
Given: the sum of their reciprocals = 3/5
In equation form this is:
1/x + 1/(5x) = 3/5
Multiply through by x:
1 + 1/5 = (3/5)x -> (3/5)x = 6/5 -> x = (5/3)(6/5) = 2
So the other number = 5*2 = 10
Ans: 2, 10